### v2

@@ Abstract
Integrating multiple observational studies for meta-analysis has sparked much interest. The presented R package WMAP (Weighted Meta-Analysis with Pseudo-Population) (Guha et al., 2024) addresses a critical gap in the implementation of integrative weighting approaches for multiple observational studies and causal inferences about various groups of subjects, such as disease subtypes. The package features three weighting approaches, each representing a special case of the unified weighting framework introduced by Guha and Li (2024), which includes an extension of inverse probability weights for data integration settings. It performs meta-analysis on user-inputted datasets as follows: (i) it first estimates the propensity scores for study-group combinations, calculates subject balancing weights, and determines the effective sample size (ESS) for a user-specified weighting method; and (ii) it then estimates various features of multiple counterfactual group outcomes, such as group medians and differences in group means for the mRNA expression of eight genes. Additionally, bootstrap variability estimates are provided. Among the implemented weighting methods, we highlight the FLEXible, Optimized, and Realistic (FLEXOR) method, which is specifically designed to maximize the ESS within the unified framework. The use of the software is illustrated by simulations as well as a multi-site breast cancer study conducted in seven medical centers.
@@ 1 Introduction
Motivated by these challenges, Guha and Li (2024) extended the propensity score to the multiple propensity score (MPS). Further, they proposed a general family of pseudo-populations and balancing weights that facilitate the integrative analyses and causal inferences of diverse functionals of group-specific potential outcomes. This unified framework generalizes the aforementioned weighting methods to the meta-analysis of multiple studies with multiple groups. Specifically, as explained in Section 2, IPWs and overlap weights (Li et al., 2018; Li and Li, 2019) are extended to the integrative combined (IC) and integrative generalized overlap (IGO) weights, respectively. Furthermore, by optimizing the effective sample size (ESS) of pseudo-populations within this rich family, Guha and Li (2024) invented the so-called FLEXible, Optimized, and Realistic (FLEXOR) weights and studied the properties of these estimand-agnostic and efficient weighted estimators for quantitative, categorical, or multivariate outcomes. The substantial benefits of FLEXOR relative to extensions of existing weighting methods such as IC and IGO are demonstrated using simulated and TCGA cancer datsets in that paper. We have developed an R package, WMAP, to implement the three weighting methods (Guha et al., 2024). The approach formulated by Guha and Li (2024) fills a methodological gap by pioneering a principled, estimand-agnostic integrative causal approach capable of accommodating multiple studies with multiple groups, a structure frequently encountered in contemporary integrative research. To our knowledge, the present R package is the first implementation of such a method.
The paper is structured as follows. Section 2 reviews the weighting methodology and two-stage inference procedure of the FLEXOR approach. Section 3 conducts simulations by using the package. Section 4 presents the meta-analysis of multiple breast cancer studies as a case study, describing the general workflow, illustrating the inferential procedure and presenting biologically meaningful results. Section 6 discusses future directions.
@@ 2.2 Stage 2: Unconfounded inferences of group counterfactual outcomes
and $\psi(\mathbf{\bar{\Phi}}_{z})$ is an estimator of $\psi(\boldsymbol{\lambda}^{(z)})$. Theorem 2 of Guha and Li (2024) shows that these weighted estimators are consistent and asymptotically normal. However, since $N$ may be small, the WMAP package applies bootstrap methods to estimate standard errors.
@@
3 Simulations
We used the WMAP package to generate simulated datasets to evaluate various weighting strategies for inferring the population-level characteristics of two subject groups and compared the inferential accuracies. Specifically, we analyzed these datasets by using the functions balancing.weights() and causal.estimate(). We briefly describe the generation strategy below and present the results. Following Guha and Li (2024) and in alignment with the motivating TCGA breast cancer studies, we simulated $R=500$ independent datasets, each containing $J=7$ observational studies, $K=2$ groups, and $L=1$ (i.e., univariate) outcomes for $\tilde{N}=500$ subjects. The covariate vectors for these subjects were sampled with replacement from the $N=450$ TCGA breast cancer patients. We set $\tilde{N}=500$ subjects across two simulation scenarios, termed “high” and “low,” to reflect the relative degrees of covariate similarity or balance among the $JK=14$ study-group combinations; specifically, the low similarity scenario represented higher confounding levels. We then employed the procedure in Section 2 to meta-analyze the four studies within each artificial dataset.
As an initial step for all $500$ artificial datasets, we conducted k-means clustering on the covariates, $\mathbf{X}_{1},\ldots,\mathbf{X}_{N}$, from the TCGA datasets to identify lower-dimensional structure, grouping them into $Q=10$ clusters with centers $\mathbf{q}_{1},\ldots,\mathbf{q}_{Q}\in\mathcal{R}^{p}$ and allocated covariate counts $m_{1},\ldots,m_{Q}$. For each artificial dataset $r=1,\ldots,500$, containing $\tilde{N}$ patients, we then generated the data as follows:
  1. 1.
    Generate the group proportions in the natural population by drawing $\boldsymbol{\theta}^{(r)}\sim\mathcal{D}_{K}(\mathbf{1}_{K})$, for $K=2$ groups. Define the group-covariate relationships: $\delta^{(r)}{z}(\mathbf{x})\propto 1$ if $z=1$ and $\delta^{(r)}{z}(\mathbf{x})\propto\exp\bigl{(}\omega_{0}^{(r)}+\omega_{1}^{(r)}\sum_{t=1}^{p}x_{t}/\frac{1}{N_{0}}\sum_{i=1}^{N_{0}}\sum_{t=1}^{p}x_{irt}^{(0)}\bigr{)}$ if $z=2$. Here, $\omega_{1}^{(r)}$ takes values of 1 and 0.1 in the high and low similarity scenarios, respectively, with $\omega_{0}^{(r)}$ chosen so that the average of $\delta^{(r)}{z}(\mathbf{x}{ir}^{(0)})$ across the population matches $\theta_{z}^{(r)}$.
  2. 2.
    Covariates  For each subject $i=1,\ldots,\tilde{N}$, select their covariate vector $\tilde{\mathbf{x}}_{i}^{(r)}=(\tilde{x}_{i1}^{(r)},\ldots,\tilde{x}_{ip}^{(r)})^{\prime}$ by sampling with replacement from the $N=450$ TCGA covariate vectors.
  3. 3.
    1. (a)
      MPS  Define the group-specific study propensities as follows:
      $\log\bigl{(}\delta_{S=s\mid Z=z}(\mathbf{x})/\delta_{S=1\mid Z=z}(\mathbf{x})\bigr{)}=sz\omega^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)}/\frac{1}{\tilde{N}}\sum_{i^{\prime}=1}^{\tilde{N}}\sum_{t=1}^{p}\tilde{x}_{i^{\prime}t}^{(r)}$
      for $s=2,\ldots,J$ and $z=1,2$. Set the similarity parameter $\omega^{(r)}$ to 0.5 in the high similarity scenario and 0.05 in the low similarity scenario. Assuming the group propensity scores are the same as in the natural population, the marginal propensity score (MPS) is given by $\delta_{sz}(\mathbf{x})=\delta_{s\mid z}(\mathbf{x})\delta_{z}(\mathbf{x})$. For each patient $i=1,\ldots,\tilde{N}$, evaluate their probability vector $\boldsymbol{\delta}^{(r)}(\mathbf{x}_{i})=\bigl{(}\delta_{11}^{(r)}(\mathbf{x}_{i}),\ldots,\delta_{JK}^{(r)}(\mathbf{x}_{i})\bigr{)}$.
    2. (b)
  4. 4.
    Subject-specific observed outcomes  Generate
    $Y_{i}^{(r)}\,\Big{|}\,Z_{i}=z_{i}^{(r)}\stackrel{{\scriptstyle\text{indep}}}{{\sim}}N\bigl{(}z_{i}^{(r)}\sum_{t=1}^{p}\tilde{x}_{it}^{(r)},\tau_{r}^{2}\bigr{)},$
Next, we set aside knowledge of all simulation parameters and analyzed each artificial dataset using the Section 2 procedure for the IC, IGO, and FLEXOR pseudo-populations. Define percent ESS as the effective sample size (ESS) scaled for 100 participants. For the 500 simulated datasets, Figure 1 displays boxplots of the percent ESS for the FLEXOR, IGO, and IC pseudo-populations in the low and high similarity scenarios. As expected, all three pseudo-populations achieved considerably higher ESS in the less demanding high similarity scenario, where covariates were nearly balanced even before applying the weighting methods. In both scenarios, the IC and IGO pseudo-populations showed comparable ESS. The FLEXOR pseudo-population, however, consistently achieved substantially higher ESS across all datasets and scenarios.
We employed the Stage 2 strategy described in Section 2 to make weighted inferences about the group mean differences, $\lambda^{(1)}-\lambda^{(2)}$, of the counterfactual outcomes in the two groups. Since the estimands are defined with respect to each pseudo-population, we assessed the accuracy of each estimator by comparing it to the true value of the corresponding estimand, determined using Monte Carlo methods. Figure 6 presents the absolute biases and variances for the FLEXOR, IGO, and IC pseudo-populations across the 500 artificial datasets of both similarity scenarios. For each artificial dataset and weighting method, the two performance measures were estimated using 500 independent bootstrap samples. Generally, the IGO and IC weights showed comparable performances for these data. All three methods exhibited similar accuracies and adequate coverage in the high similarity scenario, where covariates were nearly balanced across study-group combinations. However, in the more challenging low similarity scenario, FLEXOR often delivered the best results, frequently surpassing the other methods. Notably, under assumptions such as homoscedasticity, IGO weights are theoretically optimal for the group mean differences (see Li and Li, 2019, for single studies). However, the simulation design did not meet these sufficient conditions. The findings underscore the advantages of the FLEXOR strategy, which emphasizes stabilizing the balancing weights over targeting specific estimands.
Figure 1: For the 500 simulated datasets, percentage ESS boxplots for the three pseudo-populations in the high (left panel) and low (right panel) simulation scenarios with $\tilde{N}=500$ subjects.
Figure 2: *
Absolute bias; High simulation scenario
Figure 3: *
Absolute bias; Low simulation scenario
Figure 4: *
Standard deviation; High simulation scenario
Figure 5: *
Standard deviation; Low simulation scenario
Figure 6: For the 500 simulated datasets, absolute biases (top panels) and standard deviations (bottom panels) for the group mean differences of the three pseudo-populations in the high (left panels) and low (right panels) simulation scenarios. For each artificial dataset and pseudo-population, the absolute bias and standard deviation were estimated using 500 bootstrap samples.
@@ 4 Meta-analysis of multiple (multi-site) breast cancer studies
For demonstrating the use of the WMAP package, we analyze a multi-site breast cancer study from The Cancer Genome Atlas (TCGA). The study was conducted across $J=7$ medical centers, including $N=450$ patients divided into $K=2$ groups based on breast cancer subtypes: infiltrating ductal carcinoma (IDC) and infiltrating lobular carcinoma (ILC); the dataset can be downloaded from the GDC Data Portal after registration (NCI, 2022). The dataset includes $p=30$ unbalanced covariates and the $L=8$ outcomes mimic the mRNA expression levels of targeted breast cancer genes, COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B (Christopoulos et al., 2015), in the TCGA datasets. We wish to estimate and compare the counterfactual means, standard deviations, and medians of the two groups representing disease subtypes IDC and ILC. The package facilitates the meta-analysis of the TCGA breast cancer, and other similar cancer datasets, which can help us better understand cancer oncogenesis.
@@ 4.1 Data structure
The package contains an R replication file named demoAnalyze.R to reproduce the results presented in this paper. We begin by installing and loading the WMAP package and the example dataset included in the package.
@@
4.2 Workflow of analysis
4.2.1 Stage 1: balancing.weights()
Table I: Input arguments for function balancing.weights().
Argument Short description Default
S Vector of factor levels representing the $N$ study -
memberships. Takes values in $\{1,\ldots,J\}$
Z Vector of factor levels representing the $N$
group memberships. Takes values in $\{1,\ldots,K\}$
X Covariate matrix of $N$ rows and $p$ columns -
method Pseudo-population, i.e., weighting method; -
Can be "FLEXOR", "IC", or "IGO"
seed Seed for random number generation NULL
Relevant only when method="FLEXOR"; inputs ignored otherwise
naturalGroupProp Relevant only for FLEXOR pseudo-populations: -
fixed user-specific probability vector $\boldsymbol{\theta}$
num.random Number of random starting points of $\boldsymbol{\gamma}$ 40
in the iterative procedure
Table II: Output list components of function balancing.weights().
Position Names Short description
1 wt.v $N$ empirically normalized sample weights
2 percentESS Percentage sample ESS for pseudo-population
@@ X Covariate matrix of $N$ rows and $p$ columns -
[1] "FLEXOR...estimate10"
[…]
@@
Using the example dataset included with the WMAP package, we provide a step-by-step guide to the causal estimation of different features of the group-specific counterfactual outcomes for the three weighting methods, FLEXOR, integrative combined (IC), and integrative generalized overlap (IGO). As mentioned, the $K=2$ groups of the example dataset simulate the breast cancer subtypes, IDC and ILC, and the $L=8$ outcomes corresponding to the mRNA expression levels of the targeted breast cancer genes, COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B. The goal is unconfounded, covariate-balanced inference about the group counterfactual means, standard deviations, and medians, as well as counterfactual differences in group means and ratios of group standard deviations. If necessary, load the necessary packages and data, set a random seed, and set the runtime parameters:
@@ 4.3 Discussion of results
Applying causal.estimate() to compute all the estimates along with their corresponding 95% confidence intervals, we summarize the results in Table LABEL:tab1. It appears that all three methods, i.e., FLEXOR, IC, and IGO, consistently indicated no significant difference in the counterfactual mean expression level of the COL9A3 gene between the invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC) groups. However, a notable distinction in variability was observed, with IDC exhibiting significantly greater variability in expression levels compared to ILC. This elevated variability in IDC might reflect the inherently heterogeneous nature of this subtype, which is associated with a more diverse molecular landscape and potentially distinct tumor microenvironmental influences (Wang et al., 2024). Such variability could have implications for treatment responsiveness, as tumors with higher expression variability might exhibit differential sensitivity to targeted therapies. In contrast, the genes CXCL12 and IGF1 demonstrated significantly lower counterfactual mean expression levels in IDC compared to ILC. This observation aligns with the biological roles of these genes: CXCL12, a chemokine, is integral to cell migration and tumor metastasis, while IGF1 is known to play a role in cell growth, survival, and resistance to apoptosis. The reduced expression of these genes in IDC may indicate a subtype-specific divergence in signaling pathways that govern tumor progression. For instance, the lower CXCL12 expression in IDC may suggest altered stromal-tumor interactions, potentially impacting metastasis patterns. Similarly, the diminished IGF1 levels could reflect differences in the reliance on growth factor signaling between IDC and ILC, with implications for pathway-specific therapeutic targeting (Vanden Bempt et al., 2005). Interestingly, unlike COL9A3, the variability in expression levels for CXCL12 and IGF1 did not differ significantly between the subtypes. This uniformity in variability suggests that while these genes differ in their average expression levels, the regulatory mechanisms governing their expression stability may be conserved across IDC and ILC. This could indicate that the observed differences in mean expression are primarily driven by upstream genetic or epigenetic alterations rather than stochastic fluctuations (Samani et al., 2007). The unconfounded gene expression differences detected through this meta-causal analysis, encompassing both mean and variability distinctions between IDC and ILC, hold significant clinical implications. Specifically, the elevated variability in COL9A3 expression in IDC could serve as a biomarker for identifying high-risk patients who may benefit from intensified monitoring or tailored interventions. Meanwhile, the lower mean expression levels of CXCL12 and IGF1 in IDC suggest these genes could be leveraged as diagnostic markers to distinguish between subtypes or as therapeutic targets for subtype-specific treatment strategies. Furthermore, these findings may contribute to the development of novel prognostic models that integrate both mean expression levels and variability metrics to predict patient outcomes more accurately (McCart Reed et al., 2021). Finally, we note that in nearly all cases, the FLEXOR method outperformed the IC and IGO methods by producing notably tighter confidence intervals. This superior precision highlights the practical utility of FLEXOR in genomic research, where high-confidence estimates are critical for unraveling complex biological phenomena and informing translational applications. The enhanced resolution provided by FLEXOR could facilitate the identification of subtle yet clinically meaningful expression differences, further advancing personalized oncology.
@@
$\lambda^{(1)}$ $-0.06(-0.28,0.21)$ $0.02\boldsymbol{(-0.33,0.32)}$ $-0.04\boldsymbol{(-0.34,0.33)}$
$\lambda^{(2)}$ $-0.09(-0.11,0.36)$ $-0.04\boldsymbol{(-0.24,0.39)}$ $0.00\boldsymbol{(-0.33,0.46)}$
$\sigma^{(1)}$ $0.93(0.85,1.3)$ $1.09\boldsymbol{(0.8,1.32)}$ $1.04\boldsymbol{(0.8,1.34)}$
$\sigma^{(2)}$ $0.6(0.47,0.8)$ $0.67\boldsymbol{(0.43,0.83)}$ $0.68\boldsymbol{(0.42,0.87)}$
$M^{(1)}$ $-0.11(-0.33,0.08)$ $-0.12\boldsymbol{(-0.47,0.2)}$ $-0.12\boldsymbol{(-0.49,0.22)}$
$M^{(2)}$ $-0.06(-0.2,0.49)$ $-0.06\boldsymbol{(-0.37,0.53)}$ $0.06\boldsymbol{(-0.48,0.61)}$
$\lambda^{(1)}-\lambda^{(2)}$ $0.02(-0.43,0.17)$ $0.06\boldsymbol{(-0.65,0.28)}$ $-0.04\boldsymbol{(-0.63,0.39)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.54(1.23,2.4)$ $1.62\boldsymbol{(1.14,2.62)}$ $1.54\boldsymbol{(1.12,2.85)}$
@@
$\lambda^{(1)}$ $0.01(-0.18,0.3)$ $-0.04\boldsymbol{(-0.19,0.33)}$ $-0.08\boldsymbol{(-0.26,0.39)}$
$\lambda^{(2)}$ $0.39(0.28,1.07)$ $0.44\boldsymbol{(0.27,1.12)}$ $0.48(0.28,1.06)$
$\sigma^{(1)}$ $0.98(0.8,1.13)$ $0.96\boldsymbol{(0.79,1.16)}$ $0.98\boldsymbol{(0.75,1.15)}$
$\sigma^{(2)}$ $0.82(0.57,1.13)$ $0.86\boldsymbol{(0.57,1.2)}$ $0.94\boldsymbol{(0.52,1.11)}$
$M^{(1)}$ $0.09(-0.19,0.34)$ $0.06\boldsymbol{(-0.29,0.46)}$ $-0.07\boldsymbol{(-0.32,0.57)}$
$M^{(2)}$ $0.47(0.2,1.27)$ $0.54\boldsymbol{(0.14,1.22)}$ $0.72\boldsymbol{(0.11,1.22)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.38(-1.07,-0.18)$ $-0.48\boldsymbol{(-1.18,-0.11)}$ $-0.56\boldsymbol{(-1.12,-0.11)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.19(0.81,1.72)$ $1.12(0.79,1.66)$ $1.04\boldsymbol{(0.77,1.88)}$

### v3

@@ Abstract
Integrating multiple observational studies for meta-analysis has sparked much interest. The presented R package WMAP (Weighted Meta-Analysis with Pseudo-Population) (Guha et al., 2024) addresses a critical gap in the implementation of integrative weighting approaches for multiple observational studies and causal inferences about various groups of subjects, such as disease subtypes. The package features three weighting approaches, each representing a special case of the unified weighting framework introduced by Guha and Li (2024), which includes an extension of inverse probability weights for data integration settings. It performs meta-analysis on user-inputted datasets as follows: (i) it first estimates the propensity scores for study-group combinations, calculates subject balancing weights, and determines the effective sample size (ESS) for a user-specified weighting method; and (ii) it then estimates various features of multiple counterfactual group outcomes, such as group medians and differences in group means for the mRNA expression of eight genes. Additionally, bootstrap variability estimates are provided. Among the implemented weighting methods, we highlight the FLEXible, Optimized, and Realistic (FLEXOR) method, which is specifically designed to maximize the ESS within the unified framework. The use of the software is illustrated through simulations and a multi-site breast cancer case study based on a simulated dataset modeled after real TCGA data from seven medical centers.
@@ 1 Introduction
Motivated by these challenges, Guha and Li (2024) extended the propensity score to the multiple propensity score (MPS). Further, they proposed a general family of pseudo-populations and balancing weights that facilitate the integrative analyses and causal inferences of diverse functionals of group-specific potential outcomes. This unified framework generalizes the aforementioned weighting methods to the meta-analysis of multiple studies with multiple groups. Specifically, as explained in Section 2, IPWs and overlap weights (Li et al., 2018; Li and Li, 2019) are extended to the integrative combined (IC) and integrative generalized overlap (IGO) weights, respectively. Furthermore, by optimizing the effective sample size (ESS) of pseudo-populations within this rich family, Guha and Li (2024) invented the so-called FLEXible, Optimized, and Realistic (FLEXOR) weights and studied the properties of these estimand-agnostic and efficient weighted estimators for quantitative, categorical, or multivariate outcomes. The substantial benefits of FLEXOR relative to extensions of existing weighting methods such as IC and IGO are demonstrated using simulated and TCGA cancer datsets in that paper. The approach formulated by Guha and Li (2024) fills a methodological gap by pioneering a principled, estimand-agnostic integrative causal approach capable of accommodating multiple studies with multiple groups, a structure frequently encountered in contemporary integrative research. We have developed an R package, WMAP (Guha et al., 2024), to implement the three weighting methods. To our knowledge, the present R package is the first implementation of such a method.
The paper is structured as follows. Section 2 reviews the weighting methodology and two-stage inference procedure of the FLEXOR approach. Section 3 demonstrates the use of the WMAP package by analyzing the included demo dataset, which mimics the structure of a multi-site breast cancer study from The Cancer Genome Atlas (TCGA). It presents the meta-analysis of retrospective cohorts as a case study, describes the general workflow, illustrates the inferential procedure, and highlights biologically meaningful results. Section 4 conducts simulation studies using the package. An R script named Guha-Xu-Priyam-Li.R is included with the submission to replicate the results presented in the subsequent sections. Section 5 discusses future directions.
@@ 2.2 Stage 2: Unconfounded inferences of group counterfactual outcomes
and $\psi(\mathbf{\bar{\Phi}}_{z})$ is an estimator of $\psi(\boldsymbol{\lambda}^{(z)})$. Theorem 2 of Guha and Li (2024) shows that these weighted estimators are consistent and asymptotically normal. However, since $N$ may not be sufficiently large to justify asymptotic approximations, the WMAP package applies bootstrap methods to estimate standard errors.
@@ 3 Meta-analysis of multiple (multi-site) breast cancer studies
To demonstrate the use of the WMAP package, we analyze the built-in demo dataset, which contains simulated data designed to mirror the structure and characteristics of a multi-site breast cancer study from The Cancer Genome Atlas (TCGA). This illustrative dataset mimics the layout and variable types of the original TCGA study, which involved $J=7$ medical centers and $N=450$ patients, divided into $K=2$ groups based on breast cancer subtypes: infiltrating ductal carcinoma (IDC) and infiltrating lobular carcinoma (ILC). The original dataset, available from the GDC Data Portal upon registration (NCI, 2022), includes $p=30$ unbalanced covariates and $L=8$ outcomes representing mRNA expression levels for selected genes (COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B) known to be relevant in breast cancer research (Christopoulos et al., 2015). Using the demo dataset, we estimate and compare the counterfactual means, standard deviations, and medians between the IDC and ILC groups. The WMAP package thus provides a convenient and accessible tool for exploring causal meta-analysis methods on synthetic datasets that closely resemble real-world cancer studies, even without direct access to the full TCGA resource.
@@ 3.1 Data structure
We begin by installing and loading the WMAP package and the example dataset included in the package.
@@
Table I: Input arguments for function balancing.weights().
Argument Short description Default
S Vector of factor levels representing the $N$ study -
memberships. Takes values in $\{1,\ldots,J\}$
Z Vector of factor levels representing the $N$
group memberships. Takes values in $\{1,\ldots,K\}$
X Covariate matrix of $N$ rows and $p$ columns -
method Pseudo-population, i.e., weighting method; -
Can be "FLEXOR", "IC", or "IGO"
seed Seed for random number generation NULL
Relevant only when method="FLEXOR"; inputs ignored otherwise
naturalGroupProp Relevant only for FLEXOR pseudo-populations: -
fixed user-specific probability vector $\boldsymbol{\theta}$
num.random Number of random starting points of $\boldsymbol{\gamma}$ 40
in the iterative procedure
Table II: Output list components of function balancing.weights().
Position Names Short description
1 wt.v $N$ empirically normalized sample weights
2 percentESS Percentage sample ESS for pseudo-population
3.2 Workflow of analysis
3.2.1 Stage 1: balancing.weights()
@@ 3.2 Workflow of analysis
FLEXOR... estimate 10
FLEXOR... estimate 20
The output output1 is a result S3 list object of class ‘balancing_weights’, which contains:
  • wt.v: the weights for each individual.
@@
Using the example dataset included with the WMAP package, we provide a step-by-step guide to the causal estimation of different features of the group-specific counterfactual outcomes for the three weighting methods, FLEXOR, integrative combined (IC), and integrative generalized overlap (IGO). As mentioned, the $K=2$ groups of the example dataset simulate the breast cancer subtypes, IDC and ILC, and the $L=8$ outcomes corresponding to the mRNA expression levels of the targeted breast cancer genes, COL9A3, CXCL12, IGF1, ITGA11, IVL, LEF1, PRB2, and SMR3B. The goal is unconfounded, covariate-balanced inference about the group counterfactual means, standard deviations, and medians, as well as counterfactual differences in group means and ratios of group standard deviations.
If necessary, load the necessary packages and data, set a random seed, and set the runtime parameters:
@@ 3.3 Discussion of results
We used causal.estimate() to compute point estimates and 95% confidence intervals, and we summarize the results in Table LABEL:tab1. The analysis was conducted on a simulated dataset modeled after a multi-site TCGA breast cancer study. While the dataset is not based on actual patient data, it reflects key structural and biological features commonly found in real-world genomic studies. This setup allows us to assess the performance of the methods under controlled conditions.
In this simulated setting, all three methods (FLEXOR, IC, and IGO) consistently indicated no significant difference in the counterfactual mean expression of the COL9A3 gene between the invasive ductal carcinoma (IDC) and invasive lobular carcinoma (ILC) groups. However, FLEXOR identified significantly greater variability in expression for IDC, which is consistent with the known biological heterogeneity of this subtype in real datasets (Wang et al., 2024). This kind of variability could be important for understanding differential treatment responses or disease progression.
The genes CXCL12 and IGF1 showed lower counterfactual mean expression in IDC compared to ILC. These findings align with the biological functions of these genes; CXCL12 is involved in cell migration and metastasis, while IGF1 plays a role in cell growth and survival. Although these patterns were observed in simulated data, they demonstrate how meta-causal analysis can help detect group-level trends in gene expression that may be relevant in real studies (Vanden Bempt et al., 2005).
For both CXCL12 and IGF1, the variability in expression did not differ meaningfully between subtypes. This suggests that while the mean expression levels diverged, the underlying mechanisms driving variability may be similar. In actual biological systems, this could reflect regulation through shared pathways or constraints on gene expression stability (Samani et al., 2007).
Although the results are based on simulated data, they highlight how the methods in the WMAP package can support causal interpretation of group differences in gene expression, including both average levels and variability. These types of analyses may have practical applications in identifying subtype-specific markers or informing personalized treatment strategies (McCart Reed et al., 2021).
In nearly all scenarios, FLEXOR produced narrower confidence intervals than both IC and IGO. This higher precision illustrates FLEXOR’s strength in practice, particularly when theoretical conditions do not fully hold. By emphasizing stability and practical feasibility in weight estimation, FLEXOR provides a reliable approach to causal meta-analysis in genomic settings.
@@
$\lambda^{(1)}$ $-0.06(-0.23,0.21)$ $0.06\boldsymbol{(-0.31,0.32)}$ $-0.07\boldsymbol{(-0.37,0.36)}$
$\lambda^{(2)}$ $-0.09(-0.22,0.33)$ $-0.03\boldsymbol{(-0.23,0.45)}$ $-0.05\boldsymbol{(-0.25,0.52)}$
$\sigma^{(1)}$ $0.93(0.86,1.25)$ $1.13\boldsymbol{(0.8,1.27)}$ $1.04\boldsymbol{(0.78,1.34)}$
$\sigma^{(2)}$ $0.6(0.46,0.81)$ $0.69\boldsymbol{(0.45,0.86)}$ $0.68\boldsymbol{(0.43,0.86)}$
$M^{(1)}$ $-0.11(-0.31,0.08)$ $-0.12\boldsymbol{(-0.47,0.27)}$ $-0.18\boldsymbol{(-0.48,0.24)}$
$M^{(2)}$ $-0.06(-0.23,0.43)$ $0.06\boldsymbol{(-0.45,0.53)}$ $-0.06\boldsymbol{(-0.42,0.67)}$
$\lambda^{(1)}-\lambda^{(2)}$ $0.02(-0.42,0.17)$ $0.09\boldsymbol{(-0.61,0.42)}$ $-0.01\boldsymbol{(-0.59,0.35)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.54(1.17,2.31)$ $1.64\boldsymbol{(1.17,2.33)}$ $1.52\boldsymbol{(1.07,2.58)}$
@@
$\lambda^{(1)}$ $0.01(-0.15,0.27)$ $-0.11\boldsymbol{(-0.25,0.34)}$ $-0.11\boldsymbol{(-0.26,0.35)}$
$\lambda^{(2)}$ $0.39(0.29,0.97)$ $0.52\boldsymbol{(0.2,1.11)}$ $0.55\boldsymbol{(0.3,1.11)}$
$\sigma^{(1)}$ $0.98(0.81,1.11)$ $1.07\boldsymbol{(0.78,1.13)}$ $1\boldsymbol{(0.76,1.2)}$
$\sigma^{(2)}$ $0.82(0.6,1.12)$ $0.81(0.57,1.09)$ $0.83\boldsymbol{(0.51,1.13)}$
$M^{(1)}$ $0.09(-0.12,0.32)$ $0.04\boldsymbol{(-0.31,0.56)}$ $-0.07\boldsymbol{(-0.25,0.44)}$
$M^{(2)}$ $0.47(0.39,1.13)$ $0.68\boldsymbol{(0.07,1.36)}$ $0.72\boldsymbol{(0.16,1.38)}$
$\lambda^{(1)}-\lambda^{(2)}$ $-0.38(-0.99,-0.24)$ $-0.63\boldsymbol{(-1.15,-0.13)}$ $-0.66\boldsymbol{(-1.18,-0.24)}$
${\sigma^{(1)}}/{\sigma^{(2)}}$ $1.19(0.84,1.7)$ $1.32\boldsymbol{(0.79,1.74)}$ $1.21\boldsymbol{(0.73,1.94)}$